C=100+0,80Y, what is the savings function in the simple Keynesian model?
1. S=-100+0,20Y
2. S=100+0,20Y
3. 0,20
4. 5
1
Expert's answer
2020-04-20T09:54:52-0400

Saving is the rest of income after consumption. So we cam find saving curve:

"S=Y-C=Y-100-0.8Y=0.2Y-100."

Saving function:

"S = 0.2Y-100." So the right answer is 1.